OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(–10, 1) with slope –5. y = –5x – 49 y – 1 = –5(x + 10) y – 10 = –5(x + 1) y = –5x + 1

OpenStudy (anonymous):

y – 1 = –5(x + 10)

OpenStudy (anonymous):

line through point P(–10, 1) with slope –5 the standard point-slope form for an equation of a line y = mx + b where m is the slope and b the y-intercept. y = -5x + b Using ordered pair Pt(-10,1) to solve for b 1 = -5*-10 + b -49 = b y = -5x -49

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

thx god ;)

OpenStudy (anonymous):

Actually, y – 1 = –5(x + 10) and y = -5x -49 are the same.

OpenStudy (anonymous):

so wich one would it be?

OpenStudy (anonymous):

well if you expand y – 1 = –5(x + 10), you get y = -5x -49. I would use y = -5x -49 as the answer since it is in the standard form

OpenStudy (anonymous):

ok thx :)

OpenStudy (anonymous):

right